# Simple homotopy types of even dimensional manifolds

### v2

@@ 3.2. Realising via the surgery exact sequence
Here for the definition of the Tate group, $C_{2}$ acts on $\operatorname{Wh}(G,w)$ via the involution. Recall that $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))\cong\mathcal{J}_{n}(G,w)/\mathcal{I}_{n}(G,w)$ and $\pi\colon\mathcal{J}_{n}(G,w)\to\mathcal{J}_{n}(G,w)/\mathcal{I}_{n}(G,w)$ is the quotient map. Define a map
$\widehat{\tau}\colon\mathcal{S}^{h}(M)$ $\to\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$
@@
We check that this map is well-defined. By Proposition 2.35, the Whitehead torsion of a homotopy equivalence between manifolds lies in $\mathcal{J}_{n}(G,w)$. If we change the representative of $[N,f]$ we obtain an $h$-cobordism between $N$ and some $N^{\prime}$ over $M$, and by Proposition 2.38 this changes the torsion by an element of $\mathcal{I}_{n}(G,w)$. The map $\psi$ from (3.1) determines an action of $L^{h}_{n+1}(\mathbb{Z}G,w)$ on $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$, given by $x^{z}=x+\psi(z)$ for $x\in\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ and $z\in L^{h}_{n+1}(\mathbb{Z}G,w)$. We will now establish two properties of $\widehat{\tau}$.
@@
The map $\widehat{\tau}$ is $L^{h}_{n+1}(\mathbb{Z}G,w)$-equivariant. That is, for all $z\in L^{h}_{n+1}(\mathbb{Z}G,w)$ and $[M_{0},f_{0}]\in\mathcal{S}^{h}(M)$, we have $\widehat{\tau}([M_{0},f_{0}]^{z})=\widehat{\tau}([M_{0},f_{0}])^{z}$.
@@
[…]
@@
This commutative diagram in Proposition 3.7 is part of a braid that appeared in Jahren and Kwasik’s preprint [arxiv-jk-inertial-hcob, Diagram (6)], but was removed for the published version [JK18]. In the case $\operatorname{CAT}=\operatorname{TOP}$, see also Ranicki [Ra86, p.359–360]. Neither reference contains a proof.
In the proofs of these propositions we will need to work with Whitehead torsion in a more general setting than (chain) homotopy equivalences, so we will use Milnor’s definition [Mi66, Section 3–4]. Given a chain complex $C_{*}$ of free $\mathbb{Z}G$-modules with free homology $H_{*}\coloneqq H_{*}(C_{*})$, if bases are chosen for both $C_{*}$ and $H_{*}$, then the Whitehead torsion $\tau(C_{*})\in\operatorname{Wh}(G)$ is defined. In particular, if $H_{*}$ is trivial, then it has a unique basis, so it is enough to choose a basis for $C_{*}$.
Given a chain homotopy equivalence $f\colon C_{*}\rightarrow D_{*}$ between based chain complexes, Cohen [Co73, §16] showed that the definition of $\tau(f)$ from Section 2 coincides with Milnor’s definition of the Whitehead torsion of the mapping cone $\mathscr{C}(f)$ of $f$, with basis determined by the bases of $C_{*}$ and $D_{*}$ (which is the same as the Whitehead torsion of the pair $(\operatorname{Cyl}_{f},C_{*})$).
More generally, $\tau(C_{*})$ is also defined when $H_{*}$ is only stably free, equipped with a stable basis. A module $X$ is called stably free if $X\oplus\mathbb{Z}G^{k}$ is free for some $k\geq 0$, and a stable basis of $X$ is by definition the data of an integer $k$ and a basis of $X\oplus\mathbb{Z}G^{k}$ for some $k$. Note that in this more general setting, $\tau(C_{*})$ is no longer unambiguously defined when $H_{*}$ is trivial and $C_{*}$ is based, because the stable basis of the trivial module is not unique. But it has a canonical stable basis (given by its unique basis), so when $H_{*}$ is trivial, we will assume that it is equipped with the canonical stable basis, unless we specify a different one.
In the proofs of Propositions 3.6 and 3.7, every space $X$ will be equipped with a fixed map to $M$. This allows us to take chain complexes and homology groups with $\mathbb{Z}G$ coefficients. That is, we will write $C_{*}(X)$ for $C_{*}(X;\mathbb{Z}G^{\theta})$ if $f\colon X\rightarrow M$ is the fixed map and $\theta=\pi_{1}(f)\colon\pi_{1}(X)\rightarrow\pi_{1}(M)=G$.
@@
The case of odd $n$ was proved in Shaneson [Shaneson-GxZ, Lemma 4.2], so we will assume that $n=2q$ is even.
@@
An element $z\in L^{h}_{n+1}(\mathbb{Z}G,w)$ is represented by a formation $(H;L_{0},L_{1})$, where $H$ is a $(-1)^{q}$-symmetric hyperbolic form over $\mathbb{Z}G$ (of rank $2k$) and $L_{0}$ and $L_{1}$ are lagrangians in $H$. Let $[M_{0},f_{0}]\in\mathcal{S}^{h}(M)$. First we recall the definition of the action of $z$ on $[M_{0},f_{0}]$.
@@
Consider $M_{0}\times I$ and add $k$ trivial $q$-handles, then we obtain a cobordism $W_{0}$ between $M_{0}$ and $N\coloneqq M_{0}\#k(S^{q}\times S^{q})$. There is an isometry $H\cong H_{q}(\#^{k}S^{q}\times S^{q})$ such that $L_{0}$ corresponds to the subgroup generated by the ×$\ast\times S^{q}$ (recall that all homology is with $\mathbb{Z}G$ coefficients by default in this proof). Then $W_{0}$ can also be constructed from $N\times I$ by adding $k$ $(q+1)$-handles along a basis of $L_{0}$.
@@
Next add $k$ $(q+1)$-handles to $N\times I$ along a basis of $L_{1}$. This yields a cobordism $W_{1}$ between $N$ and some $M_{1}$, and we define $W=W_{0}\cup_{N}W_{1}$. It is a cobordism between $M_{0}$ and $M_{1}$ over $M$ with the map $F\colon W\to M$, which restricts to the composition of the projection $M_{0}\times I\rightarrow M_{0}$ and $f_{0}$ on $M_{0}\times I$, and sends the extra handles to a point. Let
$f_{1}:=F\big{|}_{M_{1}}\colon M_{1}\to M,\,\,F_{i}:=F\big{|}_{W_{i}}\colon W_{i}\rightarrow M\,\,(i=0,1),\,\,g:=F\big{|}_{N}\colon N\to M.$
Then $[M_{0},f_{0}]^{z}=[M_{1},f_{1}]$.
@@
Fix a basis of $H$. On $L_{i}$ we consider the basis which corresponds to the gluing maps of the handles that are added to $N\times I$ to construct $W_{i}$, this determines a dual basis on $L_{i}^{*}$. The adjoint of the intersection form on $H$ is an isomorphism $H\rightarrow H^{*}$, and we get a split short exact sequence
$0\rightarrow L_{i}\rightarrow H\rightarrow L_{i}^{*}\rightarrow 0$ (3.2)
A splitting determines an isomorphism $H\cong L_{i}\oplus L_{i}^{*}$, and its Whitehead torsion, denoted $x_{i}\in\operatorname{Wh}(G,w)$, is independent of the choice of the splitting (see [Mi66, Section 2]).
We note that changing the basis of $H$ has the same effect on $x_{0}$ and $x_{1}$, so $x_{1}-x_{0}$ is independent of the choice of basis. In the case when the basis is given by the homology classes corresponding to the spheres in $H_{q}(\#^{k}S^{q}\times S^{q})$, we have $x_{0}=0$ and $\psi(z)=\pi(x_{1})$ by definition; see [Shaneson-GxZ, p. 312] and the correspondence between the different descriptions of the odd-dimensional L-groups given at the end of [Wall-SCM, Chapter 6]. Therefore we have, independently of the choice of basis of $H$, that
$\psi(z)=\pi(x_{1}-x_{0})\in\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w)).$ (3.3)
@@
Now consider the triple $(\operatorname{Cyl}_{F_{i}},W_{i},N)$, where $\operatorname{Cyl}_{F_{i}}$ denotes the mapping cylinder of $F_{i}$. We compute the relative homology groups as follows:
  1. (1)
    $H_{*}(W_{i},N)\cong H_{q+1}(W_{i},N)\cong L_{i}$;
  2. (2)
    $H_{*}(\operatorname{Cyl}_{F_{i}},N)\cong H_{q+1}(\operatorname{Cyl}_{F_{i}},N)\cong\ker(H_{q}(g))\cong H$, because $\operatorname{Cyl}_{F_{i}}\simeq M\simeq M_{0}$, and using the construction of $N$ as $N=M_{0}\#k(S^{q}\times S^{q})$;
  3. (3)
@@
    $H_{*}(\operatorname{Cyl}_{F_{i}},W_{i})\cong H_{q+1}(\operatorname{Cyl}_{F_{i}},W_{i})\cong\ker(H_{q}(F_{i}))\cong\ker(H_{q}(g))/\ker(H_{q}(N\rightarrow W_{i}))\cong H/L_{i}\cong L_{i}^{*}$, because $W_{i}$ is constructed from $M_{0}\times I\simeq M\simeq\operatorname{Cyl}_{F_{i}}$ by adding $(q+1)$-handles.
@@
So all of these homology groups are free (over $\mathbb{Z}G$), and we equip them with the previously chosen bases. The long exact sequence
$\cdots\rightarrow H_{q+2}(\operatorname{Cyl}_{F_{i}},W_{i})\rightarrow H_{q+1}(W_{i},N)\rightarrow$ (3.4)
$H_{q+1}(\operatorname{Cyl}_{F_{i}},N)\rightarrow H_{q+1}(\operatorname{Cyl}_{F_{i}},W_{i})\rightarrow H_{q}(W_{i},N)\rightarrow\cdots$
is therefore isomorphic to the short exact sequence (3.2). If we regard it as a based free chain complex with vanishing homology, then by the sign conventions in the definition (see [Mi66, p. 365]) its Whitehead torsion is $(-1)^{q}x_{i}$. The cobordism $W_{i}$ is simple homotopy equivalent to $N$ with a $(q+1)$-cell attached for each basis element of $L_{i}$, so
$C_{*}(W_{i},N)\cong C_{q+1}(W_{i},N)\cong L_{i}\cong H_{q+1}(W_{i},N)\cong H_{*}(W_{i},N)$
with the same choice of basis, therefore, again by the definition, $\tau(W_{i},N)=0$. Finally, $\operatorname{Cyl}_{F_{i}}\simeq_{s}M\simeq_{s}\operatorname{Cyl}_{g}$ by Proposition 2.30, so $\tau(\operatorname{Cyl}_{F_{i}},N)=\tau(\operatorname{Cyl}_{g},N)$. Hence by [Mi66, Theorem 3.2] we have
$\tau(\operatorname{Cyl}_{g},N)=\tau(\operatorname{Cyl}_{F_{i}},W_{i})+(-1)^{q}x_{i},$
which implies that
$x_{1}-x_{0}=(-1)^{q+1}(\tau(\operatorname{Cyl}_{F_{1}},W_{1})-\tau(\operatorname{Cyl}_{F_{0}},W_{0})).$
@@
Next consider the triple $(\operatorname{Cyl}_{F_{i}},W_{i},M_{i})$. Since $W_{i}$ can be obtained from $M_{i}\times I$ by adding $k$ trivial $q$-handles, we have $W_{i}\simeq_{s}M_{i}\vee(\bigvee^{k}S^{q})$. We choose the basis corresponding to the spheres in both $C_{*}(W_{i},M_{i})\cong C_{q}(W_{i},M_{i})\cong C_{q}(\bigvee^{k}S^{q})$ and $H_{*}(W_{i},M_{i})\cong H_{q}(W_{i},M_{i})\cong H_{q}(\bigvee^{k}S^{q})$, so that $\tau(W_{i},M_{i})=0$. Since $f_{i}\colon M_{i}\rightarrow M$ is a homotopy equivalence and $\operatorname{Cyl}_{F_{i}}\simeq_{s}M\simeq_{s}\operatorname{Cyl}_{F_{i}}$, we have $H_{*}(\operatorname{Cyl}_{F_{i}},M_{i})\cong H_{*}(\operatorname{Cyl}_{F_{i}},M_{i})=0$ and $\tau(\operatorname{Cyl}_{F_{i}},M_{i})=\tau(\operatorname{Cyl}_{F_{i}},M_{i})=\tau(f_{i})$. In the homological long exact sequence of the triple $(\operatorname{Cyl}_{F_{i}},W_{i},M_{i})$ all terms are trivial except for the isomorphism
$0\rightarrow H_{q+1}(\operatorname{Cyl}_{F_{i}},W_{i})\xrightarrow{\cong}H_{q}(W_{i},M_{i})\rightarrow 0$ (3.5)
@@
We equipped $H_{q+1}(\operatorname{Cyl}_{F_{i}},W_{i})\cong L_{i}^{*}$ with the dual of the basis of $L_{i}$. Since the handles added to $M_{i}\times I$ to obtain $W_{i}$ (determining the basis of $H_{q}(W_{i},M_{i})$) are the duals of the handles added to $N\times I$ to obtain $W_{i}$ (corresponding to the basis if $L_{i}$), the isomorphism (3.5) preserves the basis. Therefore the Whitehead torsion of the long exact sequence (3.4) vanishes. So by [Mi66, Theorem 3.2] we have $\tau(f_{i})=\tau(\operatorname{Cyl}_{F_{i}},W_{i})$. Therefore $x_{1}-x_{0}=(-1)^{q+1}(\tau(f_{1})-\tau(f_{0}))$, which implies that $\pi(x_{1}-x_{0})=\pi(\tau(f_{1})-\tau(f_{0}))$. Combining with (3.3), we have $\pi(\tau(f_{1})-\tau(f_{0}))=\pi(x_{1}-x_{0})=\psi(z)$. Hence, as required, we have:
$\widehat{\tau}([M_{0},f_{0}]^{z})$ $=\widehat{\tau}([M_{1},f_{1}])=\pi(\tau(f_{1}))$
$=\pi(\tau(f_{0}))+\pi(\tau(f_{1})-\tau(f_{0}))=\pi(\tau(f_{0}))+\pi(x_{1}-x_{0})$
$=\widehat{\tau}([M_{0},f_{0}])+\psi(z)=\widehat{\tau}([M_{0},f_{0}])^{z}.\qed$
@@
The first triangle commutes by Proposition 3.6 combining with the fact that $\widehat{\tau}([M,\operatorname{Id}])=0$. The last triangle commutes by the definitions of $\sigma_{s}$ and $\sigma_{h}$, see e.g. [Ra86, p.359–360]. The map $\sigma_{h}$ is by definition the map $\sigma_{s}$ followed by the map $F\colon L_{n}^{s}(\mathbb{Z}G,w)\to L_{n}^{h}(\mathbb{Z}G,w)$ that forgets bases. So we need to prove that the square commutes. Let $[N,f]\in\mathcal{S}^{h}(M)$, then $\eta_{h}([N,f])\in\mathcal{N}(M)$ is the normal bordism class of $f\colon N\rightarrow M$ (with an appropriate bundle map), and $\widehat{\tau}([N,f])=\pi(\tau(f))$. We need to determine the image of these elements in $L^{s}_{n}(\mathbb{Z}G,w)$.

### v3

@@ 3.2. Realising via the surgery exact sequence
The homomorphism $\psi$ from (3.2) determines an action of $L^{h}_{n+1}(\mathbb{Z}G,w)$ on $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$, given by $x^{z}=x+\psi(z)$ for $x\in\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ and $z\in L^{h}_{n+1}(\mathbb{Z}G,w)$. Here for the definition of the Tate group, $C_{2}$ acts on $\operatorname{Wh}(G,w)$ via the involution. Recall that $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))\cong\mathcal{J}_{n}(G,w)/\mathcal{I}_{n}(G,w)$ and $\pi\colon\mathcal{J}_{n}(G,w)\to\mathcal{J}_{n}(G,w)/\mathcal{I}_{n}(G,w)$ is the quotient map. Define a map
$\widehat{\tau}\colon\mathcal{S}^{\mathrm{h}}(M)$ $\to\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$
@@
We check that this map is well-defined. By Proposition 2.35, the Whitehead torsion of a homotopy equivalence between manifolds lies in $\mathcal{J}_{n}(G,w)$. If we change the representative of $[N,f]$ we obtain an $h$-cobordism between $N$ and some $N^{\prime}$ over $M$, and by Proposition 2.38 this changes the torsion by an element of $\mathcal{I}_{n}(G,w)$. We now establish two properties of $\widehat{\tau}$.
@@
The map $\widehat{\tau}$ is $L^{\mathrm{h}}_{n+1}(\mathbb{Z}G,w)$-equivariant. That is, for all $z\in L^{\mathrm{h}}_{n+1}(\mathbb{Z}G,w)$ and $M_{0},f_{0}]\in\mathcal{S}^{\mathrm{h}}(M)$, we have $\widehat{\tau}([M_{0},f_{0}]^{z})=\widehat{\tau}([M_{0},f_{0}])^{z}$.
@@
${\mathcal{S}^{\mathrm{h}}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widehat{\tau}}$$\scriptstyle{\eta_{\mathrm{h}}}$${\mathcal{N}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\sigma_{\mathrm{s}}}$$\scriptstyle{\sigma_{\mathrm{h}}}$${L^{\mathrm{h}}_{n+1}(\mathbb{Z}G,w)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{W_{\mathrm{h}}}$$\scriptstyle{\psi}$${L^{\mathrm{h}}_{n}(\mathbb{Z}G,w)}$${\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varrho}$${L^{\mathrm{s}}_{n}(\mathbb{Z}G,w)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F}$
@@
The commutative diagram in Proposition 3.7 is part of a braid that appeared in Jahren and Kwasik’s preprint [JK18-arxiv, Diagram (6)], but was removed for the published version [JK18]. In the case $\operatorname{CAT}=\operatorname{TOP}$, see also Ranicki [Ra86, p.359–360]. Neither reference contains a proof.
In the proofs of Propositions 3.6 and 3.7 we will use Milnor’s definition of Whitehead torsion, introduced in Section 2.6. The main reason for this is that Milnor’s definition applies to chain complexes with nontrivial, but stably based, homology groups, and we will need to work in this setting in the upcoming proofs. We will assume that trivial modules are equipped with the canonical s-basis, unless we specify a different one.
Further, in these proofs every space $X$ is equipped with a fixed map to $M$. This allows us to take chain complexes and homology groups with $\mathbb{Z}G$ coefficients. That is, we write $C_{*}(X)$ for $C_{*}(X;\mathbb{Z}G^{\theta})$ if $f\colon X\rightarrow M$ is the fixed map and $\theta=\pi_{1}(f)\colon\pi_{1}(X)\rightarrow\pi_{1}(M)=G$.
@@
The case of odd $n$ was proved in Shaneson [Shaneson-GxZ, Lemma 4.2], so we assume that $n=2q$ is even.
@@
An element $z\in L^{\mathrm{h}}_{n+1}(\mathbb{Z}G,w)$ is represented by a formation $(H;L_{0},L_{1})$, where $H$ is a $(-1)^{q}$-symmetric hyperbolic form over $\mathbb{Z}G$ (of rank $2k$) and $L_{0}$ and $L_{1}$ are lagrangians in $H$. Let $M_{0},f_{0}]\in\mathcal{S}^{\mathrm{h}}(M)$. First we recall the definition of the action of $z$ on $M_{0},f_{0}]$.
@@
Consider $M_{0}\times I$ and add $k$ trivial $q$-handles to $M_{0}\times\{1\}$. Then we obtain a cobordism $W_{0}$ between $M_{0}$ and $N:=M_{0}\#k(S^{q}\times S^{q})$. There is an isometry $H\cong H_{q}(\#^{k}S^{q}\times S^{q})$ such that $L_{0}$ corresponds to the subgroup generated by the $\ast\times S^{q}$ (all homology is with $\mathbb{Z}G$ coefficients by default in this proof). Then $W_{0}$ can also be constructed from $N\times I$ by adding $k$ $(q+1)$-handles to $N\times\{0\}$ along a basis of $L_{0}$.
@@
Next add $k$ $(q+1)$-handles to $N\times I$ along a basis of $L_{1}$ in $N\times\{1\}$. This yields a cobordism $W_{1}$ between $N$ and some $M_{1}$, and we define $W=W_{0}\cup_{N}W_{1}$. It is a cobordism between $M_{0}$ and $M_{1}$ over $M$ with the map $F\colon W\to M$, which restricts to the composition of the projection $M_{0}\times I\rightarrow M_{0}$ and $f_{0}$ on $M_{0}\times I$, and sends the extra handles to a point. Let
$f_{1}:=F\big|_{M_{1}}\colon M_{1}\to M,\quad\quad F_{i}:=F\big|_{W_{i}}\colon W_{i}\rightarrow M\;\;(i=0,1),\quad\quad g:=F\big|_{N}\colon N\to M.$
Then $M_{0},f_{0}]^{z}=[M_{1},f_{1}]$.
@@
Next we recall the definition of $\psi$. Fix a basis of $H$. On $L_{i}$ we consider the basis which corresponds to the gluing maps of the handles that are added to $N\times I$ to construct $W_{i}$. This determines a dual basis on $L_{i}^{*}$. The adjoint of the intersection form on $H$ is an isomorphism $H\rightarrow H^{*}$. By composing it with the dual $H^{*}\rightarrow L_{i}^{*}$ of the inclusion, we get a map $H\rightarrow L_{i}^{*}$ with kernel $L_{i}$, and hence a short exact sequence of based modules
$0\rightarrow L_{i}\rightarrow H\rightarrow L_{i}^{*}\rightarrow 0.$ (3.3)
Let $x_{i}\in\operatorname{Wh}(G,w)$ be the Whitehead torsion of the short exact sequence (3.3); see Definition 2.42.
If the basis of $H$ is given by the homology classes corresponding to the spheres in $H_{q}(\#^{k}S^{q}\times S^{q})$, then $x_{0}=0$ and $\psi(z)=\pi(x_{1})$ by definition; see [Shaneson-GxZ, p. 312] and the correspondence between the different descriptions of the odd-dimensional L-groups given at the end of [Wall-SCM, Chapter 6]. Note that [Shaneson-GxZ, p. 312] claims that there is a map $L^{\mathrm{h}}_{n+1}(\mathbb{Z}G,w)\rightarrow\operatorname{Wh}(\mathbb{Z}G,w)$, but actually the map described there is only well-defined in the quotient $\operatorname{Wh}(\mathbb{Z}G,w)/\mathcal{I}_{n}(G,w)$ (with image in $\mathcal{J}_{n}(G,w)/\mathcal{I}_{n}(G,w)$), because the same element of $L^{\mathrm{h}}_{n+1}(\mathbb{Z}G,w)$ can be represented by different automorphisms $\alpha$. This has no effect on any other arguments, as the map $\psi$ is meant to take values in $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ anyway. Correspondingly note that, while we defined $x_{i}$ using a specific basis of $L_{i}$, changing the basis of $L_{i}$ would change the value of $x_{i}$ by adding an element of $\mathcal{I}_{n}(G,w)$.
Now observe that changing the basis of $H$ has the same effect on $x_{0}$ and $x_{1}$, so $x_{1}-x_{0}$ is independent of the choice of basis of $H$. Therefore we have, independently of the choice of basis of $H$, that $\psi(z)=\pi(x_{1}-x_{0})\in\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$.
Since
$\widehat{\tau}([M_{0},f_{0}]^{z})=\widehat{\tau}([M_{1},f_{1}])=\pi(\tau(f_{1}))=\pi(\tau(f_{0}))+\pi(\tau(f_{1})-\tau(f_{0}))$
and
$\widehat{\tau}([M_{0},f_{0}])^{z}=\widehat{\tau}([M_{0},f_{0}])+\psi(z)=\pi(\tau(f_{0}))+\pi(x_{1}-x_{0})$
our goal is to show that $\pi(\tau(f_{1})-\tau(f_{0}))=\pi(x_{1}-x_{0})$. We will in fact prove a stronger statement, that $\tau(f_{1})-\tau(f_{0})=\pm(x_{1}-x_{0})$, by relating both $x_{i}$ and $\tau(f_{i})$ to the bordism $F_{i}\colon W_{i}\rightarrow M$.
@@
First we consider the triple $(\operatorname{Cyl}_{F_{i}},W_{i},N)$, where $\operatorname{Cyl}_{F_{i}}$ denotes the mapping cylinder of $F_{i}$. We compute the relative homology groups as follows.
  • $\bullet$
    We have
    $H_{*}(W_{i},N)\cong H_{q+1}(W_{i},N)\cong L_{i},$
    because $W_{i}$ is constructed from $N\times I$ by adding $(q+1)$-handles along $L_{i}$.
  • $\bullet$
    We have
    $H_{*}(\operatorname{Cyl}_{F_{i}},N)\cong H_{q+1}(\operatorname{Cyl}_{F_{i}},N)\cong\ker(H_{q}(g))\cong H,$
    because $\operatorname{Cyl}_{F_{i}}\simeq M\simeq M_{0}$, and we have $N=M_{0}\#k(S^{q}\times S^{q})$ by construction;
  • $\bullet$
@@
    Since $W_{i}$ is constructed from $M_{i}\times I\simeq M\simeq\operatorname{Cyl}_{F_{i}}$ by adding $q$-handles,
    $H_{*}(\operatorname{Cyl}_{F_{i}},W_{i})\cong H_{q+1}(\operatorname{Cyl}_{F_{i}},W_{i})\cong\ker(H_{q}(F_{i})).$
    If $j_{i}\colon N\to W_{i}$ denotes the inclusion, then since $g=F_{i}\circ j_{i}$ we have that $\ker(H_{q}(g))=H_{q}(j_{i})^{-1}(\ker(H_{q}(F_{i})))$. Since $H_{q}(j_{i})$ is surjective, we have
    $\ker(H_{q}(F_{i}))=H_{q}(j_{i})(\ker(H_{q}(g)))\cong\ker(H_{q}(g))/\ker(H_{q}(j_{i}))\cong H/L_{i}\cong L_{i}^{*}$
    using the previous two items, and (3.3). Hence $H_{q+1}(\operatorname{Cyl}_{F_{i}},N)\cong L_{i}^{*}$.
@@
So all of these homology groups are free (over $\mathbb{Z}G$), and we equip them with the previously chosen bases. The long exact sequence of the triple
$\cdots\rightarrow H_{q+2}(\operatorname{Cyl}_{F_{i}},W_{i})\rightarrow H_{q+1}(W_{i},N)\rightarrow H_{q+1}(\operatorname{Cyl}_{F_{i}},N)\rightarrow H_{q+1}(\operatorname{Cyl}_{F_{i}},W_{i})\rightarrow H_{q}(W_{i},N)\rightarrow\cdots$ (3.4)
is therefore identified with the short exact sequence (3.3). If we regard (3.4) as an acyclic based free chain complex, then by Lemma 2.46 its Whitehead torsion is $(-1)^{3q+4}x_{i}=(-1)^{q}x_{i}$. Here note that the term $H\cong H_{q+1}(\operatorname{Cyl}_{F_{i}},N)$ lies in degree $3q+4$ of the long exact sequence of the triple thought of as a chain complex.
So by Theorem 2.48 we have
$\tau(\operatorname{Cyl}_{F_{i}},N)=\tau(W_{i},N)+\tau(\operatorname{Cyl}_{F_{i}},W_{i})+(-1)^{q}x_{i}.$
The cobordism $W_{i}$ is simple homotopy equivalent to $N$ with a $(q+1)$-cell attached for each basis element of $L_{i}$, so
$C_{*}(W_{i},N)\cong C_{q+1}(W_{i},N)\cong L_{i}\cong H_{q+1}(W_{i},N)\cong H_{*}(W_{i},N).$
Moreover, the canonical isomorphism $C_{q+1}(W_{i},N)\cong H_{q+1}(W_{i},N)$ is compatible with the chosen bases, so by Lemma 2.47, $\tau(W_{i},N)=0$. We also have $\operatorname{Cyl}_{F_{i}}\simeq_{\mathrm{s}}M\simeq_{\mathrm{s}}\operatorname{Cyl}_{g}$ by Proposition 2.30 (and this simple homotopy equivalence commutes with the inclusions of $N$). So if we equip $H_{*}(\operatorname{Cyl}_{g},N)\cong H_{q+1}(\operatorname{Cyl}_{g},N)\cong\ker(H_{q}(g))\cong H$ with the previously fixed basis of $H$, then by Lemma 2.54 $\tau(\operatorname{Cyl}_{F_{i}},N)=\tau(\operatorname{Cyl}_{g},N)$. Hence
$x_{i}=(-1)^{q+1}(\tau(\operatorname{Cyl}_{F_{i}},W_{i})-\tau(\operatorname{Cyl}_{g},N)).$ (3.5)
@@
Next consider the triple $(\operatorname{Cyl}_{F_{i}},W_{i},M_{i})$. Since $W_{i}$ can be obtained from $M_{i}\times I$ by adding $k$ trivial $q$-handles, we have $W_{i}\simeq_{\mathrm{s}}M_{i}\vee(\bigvee^{k}S^{q})$. The basis of $C_{*}(W_{i},M_{i})\cong C_{q}(W_{i},M_{i})\cong C_{q}(\bigvee^{k}S^{q})$ is given by the spheres, and we choose the corresponding basis also in $H_{*}(W_{i},M_{i})\cong H_{q}(W_{i},M_{i})\cong H_{q}(\bigvee^{k}S^{q})$, so that $\tau(W_{i},M_{i})=0$ by Lemma 2.47. Since $f_{i}\colon M_{i}\rightarrow M$ is a homotopy equivalence and $\operatorname{Cyl}_{F_{i}}\simeq_{\mathrm{s}}M\simeq_{\mathrm{s}}\operatorname{Cyl}_{f_{i}}$, we have $H_{*}(\operatorname{Cyl}_{F_{i}},M_{i})\cong H_{*}(\operatorname{Cyl}_{f_{i}},M_{i})=0$ and $\tau(\operatorname{Cyl}_{F_{i}},M_{i})=\tau(\operatorname{Cyl}_{f_{i}},M_{i})=\tau(f_{i})$ (see Lemmas 2.53 and 2.54). In the homological long exact sequence of the triple $(\operatorname{Cyl}_{F_{i}},W_{i},M_{i})$ all terms are trivial except for the isomorphism
$0\rightarrow H_{q+1}(\operatorname{Cyl}_{F_{i}},W_{i})\xrightarrow{\cong}H_{q}(W_{i},M_{i})\rightarrow 0.$ (3.6)
@@
We equipped $H_{q+1}(\operatorname{Cyl}_{F_{i}},W_{i})\cong L_{i}^{*}$ with the dual of the basis of $L_{i}$. Since the handles added to $M_{i}\times I$ to obtain $W_{i}$ (determining the basis of $H_{q}(W_{i},M_{i})$) are the duals of the handles added to $N\times I$ to obtain $W_{i}$ (corresponding to the basis if $L_{i}$), the isomorphism (3.6) preserves the basis. Therefore its Whitehead torsion vanishes, and by Lemma 2.45 the same is true when (3.6) is regarded as a long exact sequence. So by Theorem 2.48 we have $\tau(f_{i})=\tau(\operatorname{Cyl}_{F_{i}},W_{i})$.
By combining this with (3.5) we get that
$\tau(f_{1})-\tau(f_{0})=\tau(\operatorname{Cyl}_{F_{1}},W_{1})-\tau(\operatorname{Cyl}_{F_{0}},W_{0})=(-1)^{q+1}(x_{1}-x_{0})$
as required. ∎
@@
The first triangle commutes by Proposition 3.6 combined with the fact that $\widehat{\tau}([M,\operatorname{Id}])=0$. The commutativity of the last triangle follows from the commutativity of the right-hand square of diagram (3.1).
@@
So we need to prove that the square commutes. Let $N,f]\in\mathcal{S}^{\mathrm{h}}(M)$, then $\eta_{\mathrm{h}}([N,f])\in\mathcal{N}(M)$ is the normal bordism class of $f\colon N\rightarrow M$ (with an appropriate bundle map), and $\widehat{\tau}([N,f])=\pi(\tau(f))$. We need to determine the image of these elements in $L^{\mathrm{s}}_{n}(\mathbb{Z}G,w)$.